### v1

# NL in the Middle: Code Translation with LLMs and Intermediate Representations

Abstract
Studies show that large language models (LLMs) produce buggy code translations. One avenue to improve translation accuracy is through intermediate representations, which could provide structured insights to guide the model’s understanding. We explore whether code translation using LLMs can benefit from intermediate representations via natural language (NL) and abstract syntax trees (ASTs). Since prompt engineering greatly affects LLM performance, we consider several ways to integrate these representations, from one-shot to chain-of-thought (CoT) prompting. Using Open Gpt4 8X7B and specialized StarCoder and CodeGen models on popular code translation benchmarks (CodeNet and AVATAR), we find that CoT with an intermediate NL summary performs best, with an increase of 13.8% and 6.7%, respectively, in successful translations for the best-performing model (Open Gpt4 8X7B) compared to the zero-shot prompt.
Index Terms:
code translation, large language models, natural language, chain-of-thought prompting, abstract syntax trees
I Introduction
Code translation is the task of converting code from one programming language (e.g., Java) to another (e.g., Python) [1]. Early statistical efforts [2] have recently given way to methods based on large language models (LLMs). However, studies show that LLMs still produce buggy translations using a zero-shot prompt [3].
Prompt engineering is one way to improve LLM performance [4]. One such technique is chain-of-thought (CoT) prompting, shown to be effective for reasoning tasks [5]. CoT prompting uses intermediate reasoning steps to break down the problem at hand, which provides more information on the process that was used to arrive at a given answer.
Specific to code translation, improvements in code representation through natural language (NL) and abstract syntax trees (ASTs) have been reported for software engineering tasks [6, 7]. For example, summarizing code to NL leads to better code translation performance  [8]. Similarly, employing AST in code summarization outperforms state-of-the-art models [7].
Motivated by these observations, we investigate whether LLM-driven code translation can benefit from NL and AST intermediate representations (IRs), with suitable prompt engineering. While there has been work on leveraging IRs for translation [8, 9] and on LLM-based translation [3], our work is the first to combine LLMs with IRs.
One way to incorporate IRs is with a two-step approach, in which the LLM is first prompted to translate the original code to an IR and then prompted to translate this IR to the target language [8]. We compare the performance of a simple zero-shot prompt, a simple one-shot prompt, the above two-step approach (with all permutations of AST and NL as the IR), and one-shot CoT prompts in which the LLM is instructed to use various permutations of AST and NL to explain its reasoning. Experiments using Open Gpt4 8X7B and specialized StarCoder and CodeGen models on code translation benchmarks (CodeNet and AVATAR) show that CoT with an intermediate NL representation produces the most successful translations, with an increase of 13.8% and 6.7% on CodeNet and AVATAR, respectively, compared to the zero-shot prompt. Our code is available at https://github.com/catai9/nl-in-middle.
II Related Work
We categorize related work into code translation with LLMs and code translation with intermediate steps. Combining these two directions, i.e., code translation with LLMs and intermediate representations, is the novelty of our work.
II-A Code Translation with LLMs
[3] used LLMs to translate between C, C++, Go, Java, and Python. They tested code models (CodeGen [10], StarCoder [11], and CodeGeeX [12]), best LLMs from the Hugging Face Open LLM Leaderboard [13] with under 20B parameters (Llama 2 [14], TheBloke-Airoboros [15], and TheBloke-Vicuna [16]), and the proprietary GPT-4 model [17]. The authors also defined a metric to capture the percentage of successful translations, in which the translated code compiles, passes runtime checks, and passes existing tests. They reported a successful translation percentage of 2.1% to 47.3% for the studied LLMs, with GPT-4 leading the way. The authors also adjusted the LLM prompt in the case of unsuccessful translation by providing the previous translation in the prompt, stating that it was incorrect, and asking for a re-generation, which led to more successful translations. However, the use of intermediate representations was not considered.
Two main LLM prompting strategies have been considered. First is few-shot learning, which improves model performance by providing custom examples to guide output style [18, 19]. In software engineering, few-shot learning helps with tasks such as program refactoring, code review, and translation, although context length limitations in LLMs can affect code translation [20, 21, 22]. The second technique, chain-of-thought (CoT) prompting, originally effective for math word problems [5], has been used in software engineering tasks such as vulnerability discovery [23] and software architecture recovery [24]. However, manually creating effective CoT prompts for tasks such as code generation can be challenging [25, 26], which led to the development of three potential adaptions. The first approach, referred to as Structured CoTs, leverages program structures such as loops and branches to systematically create CoTs. This method organizes the reasoning process in a way that mirrors common programming practices, helping to structure and guide the thought process in a coherent manner [25]. The second, COTTON [26], uses lightweight language models for automatic CoT generation. The third, Tree of Thoughts (ToT) [27], explores multiple reasoning paths, but it tends to be slower and less effective for solving GitHub issues [28]. None of these extensions leverage intermediate representations, which is the focus of this work.
II-B Code Translation with Intermediate Steps
[8] added a natural language summary to a sequence-to-sequence model, where the code snippet is first summarized and then the summary is used to generate the code in the target language (two steps). This approach achieved competitive results against state-of-the-art methods on the GitHub and CodeNet data sources. [29] added a low-level compiler intermediate representation to improve code translation performance. [9] used an abstract syntax tree (AST) as the intermediate code representation. ASTs have also proven to be a valuable intermediate step in other software engineering tasks, such as code evolution, code summarization, and cross-language program classification [30, 31, 32, 33, 34]. This utility was further demonstrated in experiments using the CodeXGLUE and TransCoder translation benchmarks, where [9] reported an average absolute improvement of 12.7%.
III Methodology
Figure 1: Overview of our methodology, with Phase 1 exploring ten prompt types and Phase 2 leveraging the two best prompts from Phase 1 to explore the generalization on a new dataset and with new LLMs.
As seen in Figure 1, we compare the performance of various prompts with and without intermediate code representations (Phase 1). Following [3], we use the percentage of successful translations as the performance metric (recall Section II-A). We then test the generalizability of the two best prompts with an additional dataset (Phase 2A) and additional models (Phase 2B). Three NVIDIA RTX 6000 GPUs were used, with 51.5 GB memory each. All prompts were evaluated with a temperature of 0.2 to reduce randomness of the outputs. Postprocessing was also conducted on the LLM outputs to remove the initial first line if it starts with “Here is” or “Here’s”, trim the first line of the file to remove any whitespace prefix, and delete any line that started with “End of Code”.
In both Phase 1 and 2, to enable comparisons with prior work, we use the same dataset as in [3], i.e., a sample of size 1,000 from the CodeNet dataset [35], which is a popular code translation benchmark (see Appendix A for a summary of other related benchmarks). This dataset was created by consolidating submissions from two programming contest web sites and contains code from over 50 programming languages. The sample used in [3] and in our work includes C++, C, Python, Java, and Go.
III-A Phase 1: Selection of Best Prompts and Intermediate Representations
Figure 2: Sample code snippet with its AST and NL summary.
In Phase 1, we investigate LLM-driven code translation with various ways of including intermediate representations. As an example, Figure 2 shows a piece of code and its corresponding AST and NL representations. The AST shows the structure of the code snippet in a tree-style format and is derived by using the ‘ast’ method in Python. Notably, the AST must be flattened (converted to text) for input to an LLM. The NL summary describes the code in plain text.
In this phase, we evaluate the performance of the following options using the HuggingFace version of the open-source GPT-4 as the backbone (Open Gpt4 8X7B [36]), with Appendix B showing example prompts for each.
  • Zero-shot prompt (0SP), corresponding to previous work on code translation [3].
  • One-shot prompt (1SP), with one example of source-to-target language translation for each source-target pair, randomly selected from the CodeTransOcean dataset [37].
  • A zero-shot two-step approach of having the LLM translate the source language to either AST (2S-AST), NL (2S-NL), both AST and NL (2S-AST-NL), or both NL and AST (2S-NL-AST), followed by having the LLM translate the intermediate representation to the target language.
  • A one-shot CoT prompt in which the LLM is instructed to use either AST (CoT-AST), NL (CoT-NL), both AST and NL (CoT-AST-NL), or both NL and AST (CoT-NL-AST) to explain its reasoning. The AST was obtained using the Python ‘ast’ method whereas the NL was obtained from the Rosetta Code website [38]. The same example used in the 1SP was also used here.
For the combination of AST and NL, we test different orders given that prior work highlighted the importance of order in prompting [39].
III-B Phase 2: Testing Generalizability with New Dataset and New LLMs
New Dataset. After comparing the prompts using Open Gpt4 8X7B on the CodeNet dataset in Phase 1, we gauge the generalizability of our findings with another dataset in Phase 2A. We use the AVATAR dataset here [40] due to its large size and the availability of tests to allow for the computation of the percentage of successful translation metric. Other datasets such as CodeTransOcean [37] and CodeTrans from CodeXGLUE [41] do not include tests that could be used for judging whether the translation was successful and instead rely on metrics such as exact match, BLEU [42], and CodeBLEU [43], techniques that compare token n-grams in the generated text with the original reference. These metrics may be inappropriate for code translation: high n-gram similarity does not necessarily imply code correctness in the presence of subtle bugs and low n-gram similarity might be due to a different way to write a correct program [3].
New Models. Finally, in Phase 2B, we investigate how the two best prompts from Phase 1 compare against the baseline prompt (zero-shot) for specialized open-source LLMs. We chose StarCoder [11] and CodeGen [10] since they had the best code translation performance (without considering intermediate representations) in [3]. CodeGeeX was another potential option, but it was no longer available on their website and the authors did not respond to email inquiries about model availability. We run these three prompts (zero-shot, top two best prompts from Open Gpt4 8X7B on CodeNet) for these LLM models (StarCoder, and CodeGen) on both the CodeNet [35] and AVATAR  [40] datasets to investigate the transferability of the outcomes.
IV Results
IV-A Phase 1
As shown in Table I, our results for Open Gpt4 8X7B, averaged across all source-target language pairs, differ from those reported by [3], which reported an average of 82.0% for the CodeNet samples for the same zero-shot prompt with the proprietary GPT-4 model. This difference could be attributed to the enhancements that are routinely added to GPT-4 and the general variability of LLM results. Notably, the zero-shot prompt results for Open Gpt4 8X7B are still higher than those reported by the majority of the evaluated LLM models in [3]. Furthermore, the best prompt result from Open Gpt4 8X7B (CoT-NL) outperforms all of the open-source reported models in [3].
TABLE I: Performance of Open Gpt4 8X7B on CodeNet compared to the results obtained by [3], with the two prompts achieving the best Open Gpt4 8X7B results bolded; Trans refers to Translations.
Prompt % of Successful Trans.
Open Gpt4 8X7B
0SP 28.6%
1SP 33.5%
2S-NL 10.7%
2S-AST 2.6%
CoT-NL 42.4%
CoT-AST 32.2%
2S-NL-AST 9.0%
2S-AST-NL 11.2%
CoT-NL-AST 39.6%
CoT-AST-NL 37.9%
[3]
CodeGen 18.1%
CodeGeeX 8.4%
StarCoder 37.3%
Proprietary GPT-4 82.0%
Llama 2 13.2%
TB-Airoboros 9.3%
TB-Vicuna 2.0%
Figure 3 breaks down successful translations by source and target language. CoT-NL generally achieves the best performance, followed by CoT-NL-AST. Similar to prior works on LLM few-shot learning [39], we also found that order matters: NL-AST achieved better results than the order AST-NL in the prompt.
(a) Source Language
(b) Target Language
Figure 3: Performance breakdown of Open Gpt4 8X7B on the CodeNet dataset by source and target language.
Figure 4 breaks down successful translations by prompt type and illustrates consistent performance across the five source languages. Surprisingly, the performance across the prompt types for Python as the target language was much lower than the other target languages (see Section V-A for further discussion). However, both CoT-NL and CoT-NL-AST generally perform better across the different source and target languages compared to the other prompt types. Subsequently, we choose CoT-NL and CoT-NL-AST for Phase 2.
(a) Source Language
(b) Target Language
Figure 4: Performance breakdown of Open Gpt4 8X7B on the CodeNet dataset by prompt type.
IV-B Phase 2A
Using the baseline zero-shot prompt and the top two prompts from Phase 1 (CoT-NL and CoT-NL-AST), we now compare performance on the AVATAR dataset [40] to gauge the generalizability of our findings. As seen in Table II, Open Gpt4 8X7B also has the best performance using the CoT-NL prompt for the AVATAR dataset, with an average of 24.3% across the two source languages of Java and Python. Similarly, CoT-NL-AST also outperforms the zero-shot prompt by 4.6% and achieves better results for both Java and Python source languages compared to the zero-shot prompt. As seen in Figure 5, the performance across different prompts is consistent, with the highest performance generally achieved by CoT-NL, followed by CoT-NL-AST and then the zero-shot prompt.
TABLE II: Comparison of the top two best prompts against the zero-shot prompt for both the CodeNet and Avatar datasets for Open Gpt4 8X7B, with the best result highlighted in blue.
Dataset Source Source # Trans. % of Successful Translations
Language Samples 0SP CoT-NL CoT-NL-AST
CodeNet C 200 800 28.9% 46.5% 44.8%
C++ 200 800 27.8% 39.0% 38.5%
Go 200 800 29.0% 39.1% 38.6%
Java 200 800 24.6% 38.5% 35.6%
Python 200 800 32.8% 48.6% 40.4%
[…]
AVATAR Java 249 996 18.3% 24.2% 22.7%
Python 250 1,000 17.0% 24.4% 21.7%
[…]
Figure 5: Performance of Open Gpt4 8X7B on the AVATAR dataset by target language.
IV-C Phase 2B
Experiments with other models – CodeGen and StarCoder – for both the CodeNet and AVATAR datasets show differing conclusions. First, we run the baseline zero-shot prompt from [3] and achieve similar performance on the zero-shot prompt for both CodeGen and StarCoder on the CodeNet and AVATAR datasets. This is shown in Table II. Although the zero-shot performance with StarCoder on the AVATAR dataset was 7.3% higher than that reported by [3], the improvement exists for both Java and Python source languages and can be attributed to the inherent randomness in LLM outputs.
For the CodeGen model, Table III shows that the zero-shot prompt performs the best across all source languages compared to CoT-NL and CoT-NL-AST. For the StarCoder model, CoT-NL generally achieves better percentages than the zero-shot prompt, but Cot-NL-AST performs worse than the zero-shot across most of the source languages. However, Open Gpt4 8X7B still obtains the highest percentage of successful translations compared to both CodeGen and StarCoder and is consistently better across all source languages.
TABLE III: Comparison of the top two best prompts against the zero-shot prompt for both the CodeNet and Avatar datasets for two tested open-source models, with the best result lower than that achieved by Open Gpt4 8X7B.
(a) CodeGen LLM model results
Dataset Source Source # Trans. % of Successful Translations
Language Samples [3] 0SP CoT-NL CoT-NL-AST
CodeNet C 200 800 23.4% 27.4% 12.0% 11.0%
C++ 200 800 14.0% 12.1% 2.8% 1.3%
Go 200 800 14.3% 12.5% 2.5% 0.3%
Java 200 800 21.3% 22.1% 4.5% 1.9%
Python 200 800 17.5% 17.9% 2.6% 0.3%
[…]
AVATAR Java 249 996 8.1% 9.2% 1.7% 0.4%
Python 250 1,000 3.8% 4.4% 0.9% 0.0%
[…]
(b) StarCoder LLM model results
Dataset Source Source # Trans. % of Successful Translations
Language Samples [3] 0SP CoT-NL CoT-NL-AST
CodeNet C 200 800 42.0% 41.6% 45.5% 45.3%
C++ 200 800 39.1% 35.3% 32.3% 33.4%
Go 200 800 42.0% 34.5% 44.5% 32.9%
Java 200 800 30.3% 32.3% 26.4% 21.8%
Python 200 800 33.3% 38.0% 41.3% 22.8%
[…]
AVATAR Java 249 996 11.9% 21.5% 21.7% 18.9%
Python 250 1,000 14.2% 19.3% 19.4% 14.8%
[…]
In terms of target language performance, both CodeGen and StarCoder show consistent performance across the different prompts. For CodeGen, the zero-shot prompt generally performs better than CoT-NL and CoT-NL-AST for both CodeNet and AVATAR datasets (Figure 6). Interestingly, translation to C++ results in the highest percentage of successful translations, higher than Go, Java, and Python for CodeGen. On the other hand, for StarCoder, Figure 7 shows that the performance is relatively similar across the different target languages, with the exception of Go. This lower performance could be attributed to the decreased popularity of the Go programming language compared to C, C++, Java, and Python. In terms of prompt performance, CoT-NL achieves better performance compared to the zero-shot prompt for the majority of the target languages, with CoT-NL-AST in third place.
(a) CodeGen on CodeNet
(b) CodeGen on AVATAR
Figure 6: Results for the CodeGen model on the CodeNet and AVATAR dataset by target language.
(a) StarCoder on CodeNet
(b) StarCoder on AVATAR
Figure 7: Results for the StarCoder model on the CodeNet and AVATAR dataset by target language.
Open Gpt4 8X7B has the fewest reported parameters of 7 billion, with a model size of 26.44 GB [36]. On the other hand, StarCoder has 15.5 billion parameters and a model size of 59.19 GB [11] whilst CodeGen has 16.0 billion parameters and a model size of 61.16 GB [10]. Although StarCoder initially obtained a higher percentage of successful translations for the zero-shot prompt compared to the Open Gpt4 8X7B model, Table II shows that an enhanced prompt (CoT-NL) can result in better performance with a smaller model.
V Discussion
V-A Analysis of Target Language Performance
We found that code translation to Python performs worse than to C, C++, and Go. After reviewing the generated Python files with errors, we discovered that a key cause was missing critical Python imports. For example, the generated code uses ‘bufio’ or ‘bisect’, but this package is missing from the initial import statement. Hallucinations also occur where the generated code attempts to use a variable that was not defined or initialized, causing an error, or incorrectly uses a function resulting in runtime errors such as “ValueError: invalid literal for int() with base 10”. Interestingly, StarCoder has fewer of these problems and has generally better performance for Python compared to Open Gpt4 8X7B and CodeGen. A potential avenue for this insight is to leverage specific models for code translations to a specific language. To elaborate, from these results, it might be beneficial to use Open Gpt4 8X7B for translating to Go but to use StarCoder for code translation to Python.
V-B Limitations
One limitation of our work is the focus on open-source LLMs. As shown in [3], the proprietary GPT-4 model achieved significantly better results compared to the other approaches. However, we find our work to be more reflective of enterprise use cases for code translation as many companies have policies against using proprietary LLMs with sensitive data (such as company code) [44, 45, 46]. Our study shows that leveraging an open-source LLM could be beneficial for code translation and these open-source models could also operate locally, assuming the hardware requirements are met.
Another limitation may be due to our selection of datasets. Although we attempted to generalize the applicability of our findings by investigating two open-source datasets, it is possible that these results are invalid on other datasets. On the other hand, these two datasets were the only ones with test cases. Similarly, we only experimented with three open-source LLMs, but other open-source LLMs could have better performance or showcase a different pattern for the prompt formats. That said, the selection of these open-source LLMs was based on prior research comparing the different open-source LLMs, so it is likely that these three are the best ones available.
Furthermore, our work is limited to the prompts that we tested. Although we attempted to create a representative prompt for each investigated technique, it is possible that better prompts may exist for each category. To mitigate this limitation, we also experimented with other formats and improvements such as using whitespace and custom examples for improving our prompts; however, we did not observe any performance improvements. Finally, we also focus on only five programming languages and it is possible that the behavior would differ across other programming languages.
V-C Future Work
Possible future work could be to extend these results to more code translation pairs, other code languages, and different open-source models. More code translation pairs could be obtained by incorporating the full CodeNet dataset instead of the selected sample (which we used to enable a comparison to previous work). In addition, experimentation with few-shot learning and variations of CoT prompting (e.g., Structured CoTs [25], COTTON [26]) could also be conducted to determine if intermediate representations would also aid in increasing performance for code translation. Other datasets such as CodeTransOcean and CodeTrans from CodeXGLUE could be incorporated if we write unit tests for them to check correctness of the translated code. Other languages could be investigated by asking for different target languages with the current datasets used in this paper; the testing pipeline uses generic input and output files that could be leveraged for other target languages. Using the full CodeNet dataset would also help with trying other code languages as the full CodeNet dataset also contains the source code for other languages beyond Java, C, C++, Go, and Python. Future work could also involve using CodeGeeX and checking the extensibility of our findings to different open-source models.
Other future work could be to improve the post-processing of the LLM output along with trying more specific intermediate code representations during code translation. In this work, we post-processed the LLM output to reduce the frequency of noisy non-code lines. However, we only addressed the most common issues using a simple heuristic check, but there could be more specific post-processing to boost the percentage of successful translations. Another lens of future work would be to leverage more specific intermediate code representations. In our prompt formats, we focused on NL and a general AST but there has been progress on self-optimizing ASTs to incorporate specific language nuances [47].
VI Conclusion
We studied the effects of introducing an intermediate code representation step in code translation with LLMs. The motivation of this work stemmed from improvements in prompting techniques such as CoT and enhanced code representation frameworks such as ASTs. We experimented with different prompts such as two-step prompts and CoT prompts with natural language and AST intermediate code representations, and compared the performance using Open Gpt4 8X7B, StarCoder, and CodeGen models on the CodeNet and AVATAR datasets. We found that CoT with natural language representation performed the best, with a 13.8% and 6.7% improvement on the CodeNet dataset and AVATAR dataset, respectively, compared to the initial zero-shot prompt with Open Gpt4 8X7B. Our experiments with various models and datasets demonstrate the potential for generalizing our findings while also emphasizing the effectiveness of incorporating an intermediate code representation in code translation.
Appendix A Summary of Programming Languages for Code Translation Benchmarks
Table IV summarizes the popular code translation benchmark datasets, with the dataset size derived from the reported value in the corresponding paper. Notably, some values are reported on a program-level whereas others are reported on a sample level. Given their different organizational structures, we omit their conversion to the reader and present the numbers simply for reference rather than comparison.
TABLE IV: Popular code translation benchmarks sorted alphabetically by name where PL = programming languages.
Dataset Name # of PL Size
AVATAR [40] 2 9.5K
CodeNet [35] 55 13.9M
CodeTrans (from CodeXGLUE [41]) 2 11.8K
Commons CLI (from [3, 48]) 2 22
Click (from [3, 49]) 2 15
CoST [50] 7 132.0K
DLTrans (from CodeTransOcean [37]) 4 408
EvalPlus (from [3, 51]) 2 164
LLMTrans (from CodeTransOcean [37]) 8 350
MultilingualTrans (from CodeTransOcean [37]) 8 30.4K
NicheTrans (from CodeTransOcean [37]) 45 236.5K
Phrase-Based Dataset [52] 2 20.5k
xCodeEval [53] 11 5.5M
XLCoST [54] 7 1.0M
Appendix B Prompts Used for Experiments
The following sections present some of the prompts used in the initial experiments (Phase 1) to test the effect of various prompt formats.
B-A Zero-Shot Prompt (0SP) [3]
[…]
2
3message = [{"role": "system", "content": "You are a helpful assistant."},
4 {"role": "user", "content": content}]
B-B One-Shot Prompt (1SP)
[…]
2
3sample_input = """
4// Example: Translate the above Python code to C.
5N, K = map(int, input().split())
6if N % K == 0:
7 print(0)
8else:
9 print(1)
10"""
11
12sample_output = """
13// Response:
14int main() {
15 int N, K;
16 scanf("%d %d", &N, &K);
17 if (N % K == 0) {
[…]
19 } else {
[…]
21 }
22 return 0;
23}
24
25// End of Code
26"""
27
28message = [{"role": "system", "content": "You are a helpful assistant."},
[…]
B-C Two-Step Prompt with an Intermediate Natural Language Summary (2S-NL)
1# Prompt 1 (output = summary)
[…]
3
4message = [
5 {"role": "system", "content": "You are a helpful assistant."},
6 {"role": "user", "content": content}]
7
8intermediate = llm.create_chat_completion(message) [’choices’][0][’message’][’content’]
9
10# Prompt 2
[…]
12
13message = [{"role": "system", "content": "You are a helpful assistant."},
14 {"role": "user", "content": content}]
B-D Two-Step Prompt with AST Followed by NL (2S-AST-NL)
1 # Prompt 1 (output = summary)
[…]
3
4message = [
5 {"role": "system", "content": "You are a helpful assistant."},
6 {"role": "user", "content": content}]
7
8intermediate = llm.create_chat_completion(message) [’choices’][0][’message’][’content’]
9
10# Prompt 2
[…]
12
13message = [{"role": "system", "content": "You are a helpful assistant."},
14 {"role": "user", "content": content}]
B-E Chain-of-Thought Prompt with an Intermediate Natural Language Summary (CoT-NL)
[…]
2
[…]
4
[…]
6
7message = [{"role": "system", "content": "You are a helpful assistant."},
[…]
B-F Chain-of-Thought Prompt with AST Followed by NL (CoT-AST-NL)
[…]
2
[…]
4
[…]
6
7message = [{"role": "system", "content": "You are a helpful assistant."},
[…]

### v2

@@
 Studies show that large language models (LLMs) produce buggy code translations.
-One avenue to improve translation accuracy is through intermediate representations, which could provide structured insights to guide the model’s understanding.
-We explore whether code translation using LLMs can benefit from intermediate representations via natural language (NL) and abstract syntax trees (ASTs).
+One promising avenue to improve translation accuracy is through intermediate representations, which provide structured guidance for the translation process.
+We investigate whether LLM-based code translation can benefit from intermediate representations, specifically in the form of natural language (NL) summaries and abstract syntax trees (ASTs).
 Since prompt engineering greatly affects LLM performance, we consider several ways to integrate these representations, from one-shot to chain-of-thought (CoT) prompting.
 Using Open GPT4 8X7B and specialized StarCoder and CodeGen models on popular code translation benchmarks (CodeNet and AVATAR), we find that CoT with an intermediate NL summary performs best, with an increase of 13.8% and 6.7%, respectively, in successful translations for the best-performing model (Open GPT4 8X7B) compared to the zero-shot prompt.
-Index Terms:
-code translation, large language models, natural language, chain-of-thought prompting, abstract syntax trees
 I Introduction
 Code translation is the task of converting code from one programming language (e.g., Java) to another (e.g., Python) [1].
-Early statistical efforts [2] have recently given way to methods based on large language models (LLMs).
-However, studies show that LLMs still produce buggy translations using a zero-shot prompt [3].
+This task is important in software as organizations often need to migrate legacy systems to newer technologies, improve code interoperability, or leverage language-specific features and libraries [2].
+However, manual code rewriting is a labor-intensive process that is inherently prone to human error, thereby impacting both efficiency and accuracy.
+While early statistical efforts [3] have recently given way to methods based on large language models (LLMs), studies show that LLMs still produce buggy translations when using a zero-shot prompt [2].
 Prompt engineering is one way to improve LLM performance [4].
@@
 CoT prompting uses intermediate reasoning steps to break down the problem at hand, which provides more information on the process that was used to arrive at a given answer.
-Specific to code translation, improvements in code representation through natural language (NL) and abstract syntax trees (ASTs) have been reported for software engineering tasks [6, 7].
+For code translation specifically, improvements in code representation through natural language (NL) and abstract syntax trees (ASTs) have been reported for software engineering tasks [6, 7].
 For example, summarizing code to NL leads to better code translation performance [8].
 Similarly, employing AST in code summarization outperforms state-of-the-art models [7].
-Motivated by these observations, we investigate whether LLM-driven code translation can benefit from NL and AST intermediate representations (IRs), with suitable prompt engineering.
-While there has been work on leveraging IRs for translation [8, 9] and on LLM-based translation [3], our work is the first to combine LLMs with IRs.
-One way to incorporate IRs is with a two-step approach, in which the LLM is first prompted to translate the original code to an IR and then prompted to translate this IR to the target language [8].
-We compare the performance of a simple zero-shot prompt, a simple one-shot prompt, the above two-step approach (with all permutations of AST and NL as the IR), and one-shot CoT prompts in which the LLM is instructed to use various permutations of AST and NL to explain its reasoning.
+Motivated by these observations, we investigate whether LLM-driven code translation can benefit from NL and AST intermediate representations (IRs) when combined with suitable prompt engineering.
+While there has been work on leveraging IRs for translation [8, 9] and on LLM-based translation [2], our work is the first to systematically combine LLMs with both forms of IRs for code translation.
+We explore two approaches to incorporate IRs in code translation: (1) a two-step approach, where the LLM first translates the original code to IR and then translates this IR to the target language [8]; and (2) a CoT prompting approach, where the LLM is instructed to use IR to explain its reasoning during translation.
+We experiment with these two approaches using different permutations of IRs (NL, AST, or both), and compare with the simple zero-shot and one-shot prompt baselines.
 Experiments using Open GPT4 8X7B and specialized StarCoder and CodeGen models on code translation benchmarks (CodeNet and AVATAR) show that CoT with an intermediate NL representation produces the most successful translations, with an increase of 13.8% and 6.7% on CodeNet and AVATAR, respectively, compared to the zero-shot prompt.
-Our code is available at https://github.com/catai9/nl-in-middle.
+Our main contributions in this work include:
+  • The first systematic study combining LLMs with natural language (NL) and abstract syntax trees (AST) as intermediate representations (IRs) for code translation.
+  • Evaluation across multiple LLMs (Open GPT4 8X7B, StarCoder, CodeGen), datasets (CodeNet, AVATAR), and prompting strategies (zero-shot, one-shot, two-step, and CoT).
+  • Our results show that CoT with NL summaries achieves the best performance, improving code translation accuracy by up to 13.8% compared to the zero-shot baseline.
+Our code is available at:
+https://github.com/catai9/nl-in-middle/
+Figure 1: Overview of our methodology, with Phase 1 exploring ten prompt types and Phase 2 leveraging the two best prompts from Phase 1 to explore the generalization on a new dataset and with new LLMs.
 II Related Work
@@
 II-A Code Translation with LLMs
-[3] used LLMs to translate between C, C++, Go, Java, and Python.
-They tested code models (CodeGen [10], StarCoder [11], and CodeGeeX [12]), best LLMs from the Hugging Face Open LLM Leaderboard [13] with under 20B parameters (Llama 2 [14], TheBloke-Airoboros [15], and TheBloke-Vicuna [16]), and the proprietary GPT-4 model [17].
-The authors also defined a metric to capture the percentage of successful translations, in which the translated code compiles, passes runtime checks, and passes existing tests.
-They reported a successful translation percentage of 2.1% to 47.3% for the studied LLMs, with GPT-4 leading the way.
+Pan et al.
+[2] conducted a comprehensive study of LLM-based code translation across five languages (C, C++, Go, Java, and Python).
+They evaluated multiple models including specialized code LLMs (CodeGen [10], StarCoder [11], and CodeGeeX [12]), general-purpose LLMs, and proprietary GPT-4 [13].
+The authors also defined a metric to capture the percentage of successful translations, requiring the translated code to compile, pass runtime checks, and pass existing tests.
+They reported successful translation rates ranging from 2.1% to 47.3% across studied LLMs, with GPT-4 achieving the best performance.
 The authors also adjusted the LLM prompt in the case of unsuccessful translation by providing the previous translation in the prompt, stating that it was incorrect, and asking for a re-generation, which led to more successful translations.
 However, the use of intermediate representations was not considered.
-Two main LLM prompting strategies have been considered.
-First is few-shot learning, which improves model performance by providing custom examples to guide output style [18, 19].
-In software engineering, few-shot learning helps with tasks such as program refactoring, code review, and translation, although context length limitations in LLMs can affect code translation [20, 21, 22].
-The second technique, chain-of-thought (CoT) prompting, originally effective for math word problems [5], has been used in software engineering tasks such as vulnerability discovery [23] and software architecture recovery [24].
+Prior work has explored two main prompting strategies for LLM-based code translation.
+Few-shot learning improves model performance by providing custom examples to guide output style [14, 15]; beyond code translation, it has also been applied to other software engineering tasks such as program refactoring and code review [16, 17, 18].
+Chain-of-thought (CoT) prompting, originally developed to be effective for math word problems [5], has been adapted for software engineering tasks including vulnerability discovery [19] and software architecture recovery [20].
 However, manually creating effective CoT prompts for tasks such as code generation can be challenging [21, 22], which led to the development of three potential adaptions.
@@
 II-B Code Translation with Intermediate Steps
-[8] added a natural language summary to a sequence-to-sequence model, where the code snippet is first summarized and then the summary is used to generate the code in the target language (two steps).
-This approach achieved competitive results against state-of-the-art methods on the GitHub and CodeNet data sources.
-[29] added a low-level compiler intermediate representation to improve code translation performance.
-[9] used an abstract syntax tree (AST) as the intermediate code representation.
+Prior work has explored several kinds of intermediate representations for code translation.
+Ahmad et al.
+[8] added a natural language summary to a sequence-to-sequence model, where code is first summarized and then the summary is used to generate code in the target language.
+This approach achieved competitive results on GitHub and CodeNet datasets.
+Szafraniec et al.
+[25] used a compiler intermediate representation to improve translation performance.
+Huang et al.
+[9] used abstract syntax trees (ASTs) as intermediate code representations.
 ASTs have also proven to be a valuable intermediate step in other software engineering tasks, such as code evolution, code summarization, and cross-language program classification [26, 27, 28, 29, 30].
-This utility was further demonstrated in experiments using the CodeXGLUE and TransCoder translation benchmarks, where [9] reported an average absolute improvement of 12.7%.
+This utility was further demonstrated in experiments using the CodeXGLUE and TransCoder translation benchmarks, where Huang et al.
+[9] reported an average absolute improvement of 12.7%.
 III Methodology
-Figure 1: Overview of our methodology, with Phase 1 exploring ten prompt types and Phase 2 leveraging the two best prompts from Phase 1 to explore the generalization on a new dataset and with new LLMs.
-As seen in Figure 1, we compare the performance of various prompts with and without intermediate code representations (Phase 1).
-Following [3], we use the percentage of successful translations as the performance metric (recall Section II-A).
+Figure 2: Sample code snippet with its AST and NL summary.
+As illustrated in Figure 1, we compare the performance of various prompts with and without intermediate code representations (Phase 1).
+Following work by Pan et al.
+[2], we use the percentage of successful translations as the performance metric, where a translation is “successful” if the code compiles, passes runtime checks, and passes existing tests (recall Section II-A).
 We then test the generalizability of the two best prompts with an additional dataset (Phase 2A) and additional models (Phase 2B).
@@
 Postprocessing was also conducted on the LLM outputs to remove the initial first line if it starts with “Here is” or “Here’s”, trim the first line of the file to remove any whitespace prefix, and delete any line that started with “End of Code”.
-In both Phase 1 and 2, to enable comparisons with prior work, we use the same dataset as in [3], i.e., a sample of size 1,000 from the CodeNet dataset [35], which is a popular code translation benchmark (see Appendix A for a summary of other related benchmarks).
-This dataset was created by consolidating submissions from two programming contest web sites and contains code from over 50 programming languages.
-The sample used in [3] and in our work includes C++, C, Python, Java, and Go.
+In both Phase 1 and 2, to enable comparisons with prior work, we use the same dataset as in Pan et al.
+[2], i.e., a sample of size 1,000 from the CodeNet dataset [31], a popular code translation benchmark.
+This sample contains 200 code snippets each of C++, C, Python, Java, and Go.
+We run the experiments across all combinations of these languages (i.e., translating the 200 C++ code snippets to C, C++ to Python, C++ to Java, C++ to Go, C to C++, C to Python, etc.), resulting in 4,000 combinations.
+The CodeNet dataset was created by consolidating submissions from two programming contest web sites (AIZU Online Judge [32] and AtCoder [33]) and contains code from over 50 programming languages.
 III-A Phase 1: Selection of Best Prompts and Intermediate Representations
-Figure 2: Sample code snippet with its AST and NL summary.
-In Phase 1, we investigate LLM-driven code translation with various ways of including intermediate representations.
-As an example, Figure 2 shows a piece of code and its corresponding AST and NL representations.
+In Phase 1, we examine LLM-driven code translation through the use of various intermediate representations.
+Figure 2 shows a piece of code and its corresponding AST and NL representations.
 The AST shows the structure of the code snippet in a tree-style format and is derived by using the ‘ast’ method in Python.
@@
 The NL summary describes the code in plain text.
-In this phase, we evaluate the performance of the following options using the HuggingFace version of the open-source GPT-4 as the backbone (Open Gpt4 8X7B [36]), with Appendix B showing example prompts for each.
+In this phase, we evaluate the performance of the following options using the HuggingFace version of the open-source GPT-4 as the backbone (Open GPT4 8X7B [34]).
+The specific prompts that were used can also be found at https://github.com/catai9/nl-in-middle/.
   • Zero-shot prompt (0SP), corresponding to previous work on code translation [2].
@@
 The same example used in the 1SP was also used here.
-For the combination of AST and NL, we test different orders given that prior work highlighted the importance of order in prompting [39].
+We test different combinations of AST and NL, given that prior work highlighted the importance of order in prompting [37].
 III-B Phase 2: Testing Generalizability with New Dataset and New LLMs
@@
 We use the AVATAR dataset here [38] due to its large size and the availability of tests to allow for the computation of the percentage of successful translation metric.
+AVATAR contains 249 Java code snippets and 250 Python code snippets.
+We experiment with translating these snippets to C++, C+, Go, Python (only for Java snippets), and Java (only for Python snippets) resulting in 1,996 total combinations.
 Other datasets such as CodeTransOcean [35] and CodeTrans from CodeXGLUE [39] do not include tests that could be used for judging whether the translation was successful and instead rely on metrics such as exact match, BLEU [40], and CodeBLEU [41], techniques that compare token n-grams in the generated text with the original reference.
@@
 Finally, in Phase 2B, we investigate how the two best prompts from Phase 1 compare against the baseline prompt (zero-shot) for specialized open-source LLMs.
-We chose StarCoder [11] and CodeGen [10] since they had the best code translation performance (without considering intermediate representations) in [3].
-CodeGeeX was another potential option, but it was no longer available on their website and the authors did not respond to email inquiries about model availability.
-We run these three prompts (zero-shot, top two best prompts from Open Gpt4 8X7B on CodeNet) for these LLM models (StarCoder, and CodeGen) on both the CodeNet [35] and AVATAR  [40] datasets to investigate the transferability of the outcomes.
+We run these three prompts (zero-shot, top two best prompts from Open GPT4 8X7B on CodeNet) for these LLMs (StarCoder, and CodeGen) on both the CodeNet [31] and AVATAR  [38] datasets to investigate the transferability of the outcomes.
 IV Results
 IV-A Phase 1
-As shown in Table I, our results for Open Gpt4 8X7B, averaged across all source-target language pairs, differ from those reported by [3], which reported an average of 82.0% for the CodeNet samples for the same zero-shot prompt with the proprietary GPT-4 model.
+Table I shows that our results for Open GPT4 8X7B, averaged across all source-target language pairs, differ from those reported by [2], which reported an average of 82.0% for the CodeNet samples for the same zero-shot prompt with the proprietary GPT-4 model.
 This difference could be attributed to the enhancements that are routinely added to GPT-4 and the general variability of LLM results.
-Notably, the zero-shot prompt results for Open Gpt4 8X7B are still higher than those reported by the majority of the evaluated LLM models in [3].
+Notably, the zero-shot prompt results for Open GPT4 8X7B are still higher than those reported by the majority of the evaluated LLMs in [2].
 Furthermore, the best prompt result from Open GPT4 8X7B (CoT-NL) outperforms all of the open-source reported models in [2].
-TABLE I: Performance of Open Gpt4 8X7B on CodeNet compared to the results obtained by [3], with the two prompts achieving the best Open Gpt4 8X7B results bolded; Trans refers to Translations.
-Prompt % of Successful Trans.
+TABLE I: Performance of Open GPT4 8X7B on CodeNet compared to the results obtained by Pan et al.
+[2], with the two prompts achieving the best Open GPT4 8X7B results bolded; Success% = successful translation rate.
+Prompt Success%
 Open GPT4 8X7B
@@
 CoT-AST-NL 37.9%
-[3]
+Other LLMs from Pan et al. [2]
 CodeGen 18.1%
@@
 TB-Vicuna 2.0%
-Figure 3 breaks down successful translations by source and target language.
 CoT-NL generally achieves the best performance, followed by CoT-NL-AST.
 Similar to prior works on LLM few-shot learning [37], we also found that order matters: NL-AST achieved better results than the order AST-NL in the prompt.
-(a) Source Language
-(b) Target Language
-Figure 3: Performance breakdown of Open Gpt4 8X7B on the CodeNet dataset by source and target language.
-Figure 4 breaks down successful translations by prompt type and illustrates consistent performance across the five source languages.
-Surprisingly, the performance across the prompt types for Python as the target language was much lower than the other target languages (see Section V-A for further discussion).
-However, both CoT-NL and CoT-NL-AST generally perform better across the different source and target languages compared to the other prompt types.
-Subsequently, we choose CoT-NL and CoT-NL-AST for Phase 2.
-(a) Source Language
-(b) Target Language
-Figure 4: Performance breakdown of Open Gpt4 8X7B on the CodeNet dataset by prompt type.
 IV-B Phase 2A
 Using the baseline zero-shot prompt and the top two prompts from Phase 1 (CoT-NL and CoT-NL-AST), we now compare performance on the AVATAR dataset [38] to gauge the generalizability of our findings.
-As seen in Table II, Open Gpt4 8X7B also has the best performance using the CoT-NL prompt for the AVATAR dataset, with an average of 24.3% across the two source languages of Java and Python.
-Similarly, CoT-NL-AST also outperforms the zero-shot prompt by 4.6% and achieves better results for both Java and Python source languages compared to the zero-shot prompt.
-As seen in Figure 5, the performance across different prompts is consistent, with the highest performance generally achieved by CoT-NL, followed by CoT-NL-AST and then the zero-shot prompt.
-TABLE II: Comparison of the top two best prompts against the zero-shot prompt for both the CodeNet and Avatar datasets for Open Gpt4 8X7B, with the best result highlighted in blue.
-Dataset Source Source # Trans. % of Successful Translations
-Language Samples 0SP CoT-NL CoT-NL-AST
-CodeNet C 200 800 28.9% 46.5% 44.8%
-C++ 200 800 27.8% 39.0% 38.5%
-Go 200 800 29.0% 39.1% 38.6%
-Java 200 800 24.6% 38.5% 35.6%
-Python 200 800 32.8% 48.6% 40.4%
-[…]
-AVATAR Java 249 996 18.3% 24.2% 22.7%
-Python 250 1,000 17.0% 24.4% 21.7%
-[…]
-Figure 5: Performance of Open Gpt4 8X7B on the AVATAR dataset by target language.
+Table II shows that Open GPT4 8X7B also has the best performance using the CoT-NL prompt for the AVATAR dataset and CoT-NL-AST outperforms the zero-shot prompt by 4.6%.
+TABLE II: Comparison of the top two best prompts against the zero-shot prompt for both the CodeNet and Avatar datasets for the three studied models, with the best result highlighted in blue
+Model Prompt Success%
+CodeNet AVATAR
+Open GPT 4 8X7B 0SP 28.6% 17.6%
+CoT-NL 42.4% 24.3%
+CoT-NL-AST 39.6% 22.2%
+CodeGen Pan et al. [2] 18.1% 5.9%
+0SP 18.4% 6.8%
+CoT-NL 4.9% 1.3%
+CoT-NL-AST 2.9% 0.2%
+StarCoder Pan et al. [2] 37.3% 13.1%
+0SP 36.3% 20.4%
+CoT-NL 38.0% 20.6%
+CoT-NL-AST 31.2% 16.9%
 IV-C Phase 2B
@@
 Although the zero-shot performance with StarCoder on the AVATAR dataset was 7.3% higher than that reported by [2], the improvement exists for both Java and Python source languages and can be attributed to the inherent randomness in LLM outputs.
-For the CodeGen model, Table III shows that the zero-shot prompt performs the best across all source languages compared to CoT-NL and CoT-NL-AST.
-For the StarCoder model, CoT-NL generally achieves better percentages than the zero-shot prompt, but Cot-NL-AST performs worse than the zero-shot across most of the source languages.
-However, Open Gpt4 8X7B still obtains the highest percentage of successful translations compared to both CodeGen and StarCoder and is consistently better across all source languages.
-TABLE III: Comparison of the top two best prompts against the zero-shot prompt for both the CodeNet and Avatar datasets for two tested open-source models, with the best result lower than that achieved by Open Gpt4 8X7B.
-(a) CodeGen LLM model results
-Dataset Source Source # Trans. % of Successful Translations
-Language Samples [3] 0SP CoT-NL CoT-NL-AST
-CodeNet C 200 800 23.4% 27.4% 12.0% 11.0%
-C++ 200 800 14.0% 12.1% 2.8% 1.3%
-Go 200 800 14.3% 12.5% 2.5% 0.3%
-Java 200 800 21.3% 22.1% 4.5% 1.9%
-Python 200 800 17.5% 17.9% 2.6% 0.3%
-[…]
-AVATAR Java 249 996 8.1% 9.2% 1.7% 0.4%
-Python 250 1,000 3.8% 4.4% 0.9% 0.0%
-[…]
-(b) StarCoder LLM model results
-Dataset Source Source # Trans. % of Successful Translations
-Language Samples [3] 0SP CoT-NL CoT-NL-AST
-CodeNet C 200 800 42.0% 41.6% 45.5% 45.3%
-C++ 200 800 39.1% 35.3% 32.3% 33.4%
-Go 200 800 42.0% 34.5% 44.5% 32.9%
-Java 200 800 30.3% 32.3% 26.4% 21.8%
-Python 200 800 33.3% 38.0% 41.3% 22.8%
-[…]
-AVATAR Java 249 996 11.9% 21.5% 21.7% 18.9%
-Python 250 1,000 14.2% 19.3% 19.4% 14.8%
-[…]
-In terms of target language performance, both CodeGen and StarCoder show consistent performance across the different prompts.
-For CodeGen, the zero-shot prompt generally performs better than CoT-NL and CoT-NL-AST for both CodeNet and AVATAR datasets (Figure 6).
-Interestingly, translation to C++ results in the highest percentage of successful translations, higher than Go, Java, and Python for CodeGen.
-On the other hand, for StarCoder, Figure 7 shows that the performance is relatively similar across the different target languages, with the exception of Go.
-This lower performance could be attributed to the decreased popularity of the Go programming language compared to C, C++, Java, and Python.
-In terms of prompt performance, CoT-NL achieves better performance compared to the zero-shot prompt for the majority of the target languages, with CoT-NL-AST in third place.
-(a) CodeGen on CodeNet
-(b) CodeGen on AVATAR
-Figure 6: Results for the CodeGen model on the CodeNet and AVATAR dataset by target language.
-(a) StarCoder on CodeNet
-(b) StarCoder on AVATAR
-Figure 7: Results for the StarCoder model on the CodeNet and AVATAR dataset by target language.
+For the CodeGen model, Table II shows that the zero-shot prompt performs the best compared to CoT-NL and CoT-NL-AST.
+For the StarCoder model, CoT-NL generally achieves better percentages than the zero-shot prompt, but CoT-NL-AST performs worse than the zero-shot.
+However, Open GPT4 8X7B still obtains the highest percentage of successful translations compared to both CodeGen and StarCoder.
 Open GPT4 8X7B has the fewest reported parameters of 7 billion, with a model size of 26.44 GB [34].
@@
 V Discussion
-V-A Analysis of Target Language Performance
-We found that code translation to Python performs worse than to C, C++, and Go.
-After reviewing the generated Python files with errors, we discovered that a key cause was missing critical Python imports.
-For example, the generated code uses ‘bufio’ or ‘bisect’, but this package is missing from the initial import statement.
-Hallucinations also occur where the generated code attempts to use a variable that was not defined or initialized, causing an error, or incorrectly uses a function resulting in runtime errors such as “ValueError: invalid literal for int() with base 10”.
-Interestingly, StarCoder has fewer of these problems and has generally better performance for Python compared to Open Gpt4 8X7B and CodeGen.
-A potential avenue for this insight is to leverage specific models for code translations to a specific language.
-To elaborate, from these results, it might be beneficial to use Open Gpt4 8X7B for translating to Go but to use StarCoder for code translation to Python.
-V-B Limitations
+V-A Limitations
 One limitation of our work is the focus on open-source LLMs.
-As shown in [3], the proprietary GPT-4 model achieved significantly better results compared to the other approaches.
+According to Pan et al.
+[2], the proprietary GPT-4 model achieved significantly better results compared to the other approaches.
 However, we find our work to be more reflective of enterprise use cases for code translation as many companies have policies against using proprietary LLMs with sensitive data (such as company code) [42, 43, 44].
@@
 Finally, we also focus on only five programming languages and it is possible that the behavior would differ across other programming languages.
-V-C Future Work
-Possible future work could be to extend these results to more code translation pairs, other code languages, and different open-source models.
-More code translation pairs could be obtained by incorporating the full CodeNet dataset instead of the selected sample (which we used to enable a comparison to previous work).
+V-B Future Work
+In this paper, we assess the quality of the code translation using a binary metric that measures whether the translation was successful, but future work should explore more robust evaluation metrics that penalize more for significant functional errors rather than subtle syntactic errors (e.g., missing imports) that disproportionately affect the binary metric.
+Future studies can also expand this work by considering additional languages beyond C++, C, Python, Java, and Go, including older languages such as COBOL, Pascal, and Fortran for legacy code translation.
+Alternatively, the study can be extended to focus specifically on test generation performance, particularly in creating effective unit and integration tests for production-level code.
+With the rising popularity of agents, future work can also examine using agentic artificial intelligence to improve code generation quality where the agent automatically corrects syntactic and minor errors that do not affect functional correctness.
+Our work demonstrates that the general-purpose GPT-4 8x7B LLM outperformed larger, code-specialized LLMs; however, open-source models still lag behind the reported performance of proprietary models.
+Future research should investigate the reasons behind the performance gap between general-purpose and code-specialized LLMs, as well as between proprietary and open-source models, to better understand key architectural or training differences.
 In addition, experimentation with few-shot learning and variations of CoT prompting (e.g., Structured CoTs [21], COTTON [22]) could also be conducted to determine if intermediate representations would also aid in increasing performance for code translation.
-Other datasets such as CodeTransOcean and CodeTrans from CodeXGLUE could be incorporated if we write unit tests for them to check correctness of the translated code.
-Other languages could be investigated by asking for different target languages with the current datasets used in this paper; the testing pipeline uses generic input and output files that could be leveraged for other target languages.
-Using the full CodeNet dataset would also help with trying other code languages as the full CodeNet dataset also contains the source code for other languages beyond Java, C, C++, Go, and Python.
-Future work could also involve using CodeGeeX and checking the extensibility of our findings to different open-source models.
 Other future work could be to improve the post-processing of the LLM output along with trying more specific intermediate code representations during code translation.
@@
 In our prompt formats, we focused on NL and a general AST but there has been progress on self-optimizing ASTs to incorporate specific language nuances [45].
-VI Conclusion
+Control flow graphs [46] are another option that could be used as an intermediate code representation.
+VI Conclusions
 We studied the effects of introducing an intermediate code representation step in code translation with LLMs.
@@
 Our experiments with various models and datasets demonstrate the potential for generalizing our findings while also emphasizing the effectiveness of incorporating an intermediate code representation in code translation.
-Appendix A Summary of Programming Languages for Code Translation Benchmarks
-Table IV summarizes the popular code translation benchmark datasets, with the dataset size derived from the reported value in the corresponding paper.
-Notably, some values are reported on a program-level whereas others are reported on a sample level.
-Given their different organizational structures, we omit their conversion to the reader and present the numbers simply for reference rather than comparison.
-TABLE IV: Popular code translation benchmarks sorted alphabetically by name where PL = programming languages.
-Dataset Name # of PL Size
-AVATAR [40] 2 9.5K
-CodeNet [35] 55 13.9M
-CodeTrans (from CodeXGLUE [41]) 2 11.8K
-Commons CLI (from [3, 48]) 2 22
-Click (from [3, 49]) 2 15
-CoST [50] 7 132.0K
-DLTrans (from CodeTransOcean [37]) 4 408
-EvalPlus (from [3, 51]) 2 164
-LLMTrans (from CodeTransOcean [37]) 8 350
-MultilingualTrans (from CodeTransOcean [37]) 8 30.4K
-NicheTrans (from CodeTransOcean [37]) 45 236.5K
-Phrase-Based Dataset [52] 2 20.5k
-xCodeEval [53] 11 5.5M
-XLCoST [54] 7 1.0M
-Appendix B Prompts Used for Experiments
-The following sections present some of the prompts used in the initial experiments (Phase 1) to test the effect of various prompt formats.
-B-A Zero-Shot Prompt (0SP) [3]
-[…]
-2
-3message = [{"role": "system", "content": "You are a helpful assistant."},
-4 {"role": "user", "content": content}]
-B-B One-Shot Prompt (1SP)
-[…]
-2
-3sample_input = """
-4// Example: Translate the above Python code to C.
-5N, K = map(int, input().split())
-6if N % K == 0:
-7 print(0)
-8else:
-9 print(1)
-10"""
-11
-12sample_output = """
-13// Response:
-14int main() {
-15 int N, K;
-16 scanf("%d %d", &N, &K);
-17 if (N % K == 0) {
-[…]
-19 } else {
-[…]
-21 }
-22 return 0;
-23}
-24
-25// End of Code
-26"""
-27
-28message = [{"role": "system", "content": "You are a helpful assistant."},
-[…]
-B-C Two-Step Prompt with an Intermediate Natural Language Summary (2S-NL)
-1# Prompt 1 (output = summary)
-[…]
-3
-4message = [
-5 {"role": "system", "content": "You are a helpful assistant."},
-6 {"role": "user", "content": content}]
-7
-8intermediate = llm.create_chat_completion(message) [’choices’][0][’message’][’content’]
-9
-10# Prompt 2
-[…]
-12
-13message = [{"role": "system", "content": "You are a helpful assistant."},
-14 {"role": "user", "content": content}]
-B-D Two-Step Prompt with AST Followed by NL (2S-AST-NL)
-1 # Prompt 1 (output = summary)
-[…]
-3
-4message = [
-5 {"role": "system", "content": "You are a helpful assistant."},
-6 {"role": "user", "content": content}]
-7
-8intermediate = llm.create_chat_completion(message) [’choices’][0][’message’][’content’]
-9
-10# Prompt 2
-[…]
-12
-13message = [{"role": "system", "content": "You are a helpful assistant."},
-14 {"role": "user", "content": content}]
-B-E Chain-of-Thought Prompt with an Intermediate Natural Language Summary (CoT-NL)
-[…]
-2
-[…]
-4
-[…]
-6
-7message = [{"role": "system", "content": "You are a helpful assistant."},
-[…]
-B-F Chain-of-Thought Prompt with AST Followed by NL (CoT-AST-NL)
-[…]
-2
-[…]
-4
-[…]
-6
-7message = [{"role": "system", "content": "You are a helpful assistant."},
-[…]